OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? A.112 B.288 C.360 D.432

OpenStudy (anonymous):

Put x on the left side. Like this, x-0.8x=72 Remember that infront of x, you have 1, so do this 1x-0.8x = 72 Now do the math there and you will get your answer.

OpenStudy (anonymous):

multiplying by 10 on RHS and LHS, 10x-8x-720=0 2x=720 x=360